Assume that today's date is February 15, 2015. Robin Hood Inc. bond is an annual-coupon bond. Par value of the bond is $1,000. How much you will pay for the bond if you purchased the bond today?

Company

Price

Coupon Rate

Maturity Date

YTM

Current Yield

Rating

Robin Hood

122.759

9.067

2-15- 2028

-

-

D
Answer #1

Coupon payment = 1000*9.067% = 90.67 Bond current price = 1000*122.759= 1227.59 Current yield = Coupon payment / Current pric